OpenStudy (anonymous):

The regular price of a child's entry ticket to a water park is $8 less than that for an adult's. The park offers half off of all entry tickets during the off-peak season. The Sandlers paid a total of $194 for 1 adult ticket and 4 child's tickets to the water park during the off-peak season. The following equation represents this situation, where x represents the regular price of an adult ticket. 194 = one-halfx + 2(x - 8) What is the regular price of a child's ticket? $68 $76 $84 $89 help?

OpenStudy (johnweldon1993):

Okay just a little change from that last one from before :) \[\frac{ 1 }{ 2 }x + 2(x - 8) = 194\] Lets distribute that 2 into the parenthesis \[\frac{ 1 }{ 2 }x + 2x - 16 = 194\] Now combine like terms \[\frac{ 5 }{ 2 }x -16 = 194\] Add 16 to both sides \[\frac{ 5 }{ 2 }x = 210\] and finally multiply both sides by 2/5 \[x = 210 \times \frac{ 2 }{ 5 }\] so what is 210 times 2? 420 and what is 420 divided by 5? 84 so it would be $84 dollars for an adults ticket....so since a child's ticket is $8 less than that we get \[84 - 8 = 76\]

OpenStudy (anonymous):

haha :) thank you! i tried finding your example but find it :) thank you :)

OpenStudy (johnweldon1993):

No problem.....and Actually in that LAST example I never finished lol.....what was at then end that I solved was the price of an adults ticket...not the childs...for the childs I had to subtract the $6 from the number that I got...oops :/ lol This one is good because I caught my mistake from before lol...but that last one is incomplete :)